There is a 5-card poker hand. How many different hands need to expect to deal between one full house and another.
Answer #1

here probability of a full house =P(selecting a denomination from 13 and then select 3 card from 4 of this denomination and than select a denomination from remaining 12 and then select 2 card from 4 of this denomination)=13C1*12C1*4C3*4C2/52C5 =0.001441

hence expected number of deal between one full house and another =1/p=1/0.001441=694.17